### v1

# Group inverses of matrices of simple symmetric directed graphs

Abstract.
A new class of simple symmetric digraphs called $\mathbb{D}$ is defined and studied here. Any digraph in $\mathbb{D}$ has the property that each non-pendant is adjacent to at least one pendant vertex. A graph theoretical description for the entries of the group inverse of a real square matrix $A$ with any digraph belonging to this class is given. We classify all the real square matrix $A$ such that the digraphs associated with $A$ and $A^{\#}$ both are in $\mathbb{D}$, that is the digraph related to $A$ is either a corona or a star digraph.
Key words and phrases:
aaa
2020 Mathematics Subject Classification:
05C12 (primary); 05C50, 05C76 (secondary)
1. Introduction
Firstly, let us recall the definition of the group inverse of a matrix, the object of primary interest here. For a real $n\times n$ matrix $A$, the group inverse, if it exists, is the matrix $X$ that satisfies the equations $AXA=A,XAX=X$ and $AX=XA$. Such an $X$ is always unique and denoted by $A^{\#}$. Any matrix $X$ that satisfies the first equation is called an inner inverse of $A$ while satisfying the second equation will be called an outer inverse. It is well-known that the group inverse of $A$ exists if and only if $rank(A)=rank(A^{2})$. For a non-singular matrix $A$, $A^{-1}=A^{\#}$. Let us recall that for a real rectangular matrix $A$, the Moore-Penrose inverse of $A$, is the unique matrix $A^{\dagger}$ that satisfies the equations $AA^{\dagger}A=A,~{}A^{\dagger}AA^{\dagger}=A^{\dagger},~{}(AA^{\dagger})^{T}=AA^{\dagger}$ and $(A^{\dagger}A)^{T}=A^{\dagger}A$. We refer the reader to [7] for more details on these notions of generalized inverses and Moore-Penrose inverses.
Now, recall some graph-theoretic notation from [15]. Let $A=(a_{ij})$ be an $n\times n$ matrix with real entries. The digraph $D(A)=(V,E)$ corresponding to $A$ is the directed graph whose vertex set is $V=\{1,2,\ldots,n\}$ and whose edge set $E$ is described by the requirement that $(i,j)\in E$ iff $a_{ij}\neq 0$. For $m\geq 1$, a sequence $(i_{1},i_{2},\ldots,i_{m},i_{m+1})$ of $m+1$ distinct vertices with edges $(i_{1},i_{2}),(i_{2},i_{3}),\ldots,(i_{m},i_{m+1})$ in $E$ is called a path of length $m$ from $i_{1}$ to $i_{m+1}$ in $D(A)$. For $m\geq 2$, a sequence $(i_{1},i_{2},\ldots,i_{m},i_{1})$ with distinct $m$ vertices $i_{1},i_{2},\ldots,i_{m},$ where $(i_{1},i_{2}),(i_{2},i_{3}),\ldots,(i_{m},i_{1})\in E,$ is called an $m$-$cycle$ (a cycle of length $m$) in $D(A)$.
A real square matrix $A=(a_{ij})$ is called combinatorially symmetric if $a_{ij}\neq 0$ iff $a_{ji}\neq 0$. Trivially, any symmetric matrix is combinatorially symmetric. Parter was the first author to use the concept of combinatorially symmetric matrices to study important problems significantly [24]. In 1974, Maybee first introduced the notion of a combinatorially symmetric matrix in [21]. Digraph $D(A)$ corresponding to a matrix $A$ is called a simple symmetric digraph if it has no loops and there is an edge in each direction between two distinct vertices i.e. the corresponding matrix $A$ is combinatorially symmetric matrix with zero diagonal. Recall that a directed graph that has a path from each vertex to every other vertex is called strongly connected. Digraph $D(A)$ corresponding to a matrix $A$ is called a tree if it is strongly connected, and all of its cycles have length 2. Clearly, a tree digraph is a simple symmetric digraph.
Suppose $A$ is a real square matrix such that $D(A)$ is a simple symmetric digraph. A cycle $(i,j,i)$ is said to be incident to $i$ as well as $j$ in $D(A)$. A vertex $i$ is called a pendant vertex if it is incident to only one 2-cycle and non-pendant vertex if it is incident to more than one 2-cycle in $D(A)$. A 2-cycle $(i,j,i)$ will be called a pendant cycle if at least one vertex $i$ or $j$ is pendant in $D(A)$, while a 2-cycle which is not pendant will be called a non-pendant cycle. A pair of vertices $i,j$ is said to be adjacent to each other if there is a 2-cycle $(i,j,i)$ in $D(A)$. For $r$ even, a set of $\frac{r}{2}$ disjoint 2-cycles in $D(A)$ given by $\{(i_{1},i_{2},i_{1}),(i_{3},i_{4},i_{3}),\ldots,(i_{r-1},i_{r},i_{r-1})\},$ is called a matching and the product $a_{i_{1},i_{2}}a_{i_{2},i_{1}}a_{i_{3},i_{4}}a_{i_{4},i_{3}}\ldots a_{i_{r-1},i_{r}}a_{i_{r},i_{r-1}}$ is called a matching product. If this set of 2-cycles has a maximum cardinality then the matching is referred to as a maximum matching and the matching product is then called a maximum matching product. The sum of all maximum matching products in $D(A)$ is denoted by $\Delta_{A}$. A matching is said to be a perfect matching if it covers all the vertices of $D(A)$. $\mathbb{M}$ and $\mathbb{M}(i)$ denote the set of all maximum matchings in $D(A)$ and the set of all maximum matchings in which vertex $i$ is matched, respectively. For a cycle $(i,j,i)$ in $D(A)$, the product $a_{ij}a_{ji}$ is called the cycle product. A sequence of $m$ 2-cycles $((i_{1},i_{2},i_{1}),(i_{2},i_{3},i_{2}),\ldots,(i_{m},i_{m+1},i_{m}))$ with $m+1$ distinct vertices $i_{1},i_{2},\ldots,i_{m+1}$ in $D(A)$ is called a cycle chain from $i_{1}$ to $i_{m+1}$ of length $m$ and denoted by $C_{m}(i_{1},i_{m+1})$. A cycle chain $C_{m}(i_{1},i_{m+1})$ is said to be an alternating cycle chain with respect to a maximum matching $M$ if cycles of $C_{m}(i_{1},i_{m+1})$ alternatively belong to $M$ and $M^{c}$, with the condition that both the first and the last cycle of $C_{m}(i_{1},i_{m+1})$ belong to $M$. For an alternating cycle chain $C_{m}(i_{1},i_{m+1})$, the product $a_{i_{1},i_{2}}a_{i_{2},i_{3}}a_{i_{3},i_{4}}\ldots a_{i_{m-1},i_{m}}a_{i_{m},i_{m+1}}$ is said to be the path product along alternating cycle chain $C_{m}(i_{1},i_{m+1})$, denoted by $P_{m}(i_{1},i_{m+1})$.
In combinatorial matrix theory, an interesting problem is giving the inverse or group inverse of a matrix using its graph structure. Here are some articles that explain how to determine the inverse [1], [4], [5], [13], [20], [25] and the group inverse [8], [12], [15], [16], [17], [23], [27]. In chemistry, the group inverse of adjacency matrices is used to determine a separation gap of a molecule by using graph energies [26]. An important parameter in quantum chemistry is the smallest positive eigenvalue. There appear to be fewer methods for computing the smallest positive eigenvalue of a matrix in contrast to a large number of techniques for computing the largest positive eigenvalue (even approximately). For a square matrix $A$, it is well known that the set of all the non-zero eigenvalues of $A^{\#}$ is the set of all reciprocals of the non-zero eigenvalues of $A$. Thus, just as is the case for nonsingular matrices, to obtain estimates for the smallest positive eigenvalue of $A$, one may consider estimates for the largest positive eigenvalue of $A^{\#}$. Also, the group inverses of matrices have applications in algebraic connectivity [19], Markov chains [11] and resistance distance [10].
We recall some well-known results on the role of generalized inverses in graph theory. In [15, Theorem 2.2], authors computed the group inverse of the bipartite matrix of the form $A=\begin{pmatrix}0&B\\ C&0\end{pmatrix}$ and presented a graphical description for the entries of $A^{\#}$ for path digraph $D(A)$. In the same article, they proposed a conjecture ([15, Conjecture 5.1]) on the entries of $A^{\#}$ when $D(A)$ is a tree digraph and this conjecture is solved recently in [23, Theorem 2.11] for a special class of trees. A block-wise description for $A^{\#}$ was given in [16, Theorem 2.1] when $B,C$ of the form $[X\ U],[Y\ V]^{T}$ respectively, with $X,Y$ non-singular and $rank(UV)=1$. Also, a graphical description for the entries of $A^{\#}$ such that $D(A)$ is a broom tree was derived in [16]. In [27], the authors gave a graphical description for entries of the adjacency matrix of arbitrary weighted trees in terms of maximum matchings and alternating paths. Group inverses matrices associated with cycle graphs were investigated in [14, 17, 26, 28].
Organization of the paper: In section 2, we define a new class $\mathbb{D}$ of simple symmetric directed graphs. We recall a block-wise description for the group inverse of a matrix $A$ with $D(A)\in\mathbb{D}$, then provide a graph theoretical description for the entries of $A^{\#}$ in Theorem 2.8. In section 3, we characterize all the real square matrix $A$ such that $D(A)$ and $D(A^{\#})$ both are in $\mathbb{D}$ in Theorem 3.5.
2. Graphical Description of $A^{\#}$ for a Matrix with Simple Symmetric Digraph
In [15], authors presented a graphical description for the entries of group inverse of a matrix $A$ with path digraph $D(A)$ and proposed a conjecture for tree digraph. Recently, in [23], we gave a formula for the entries of the group inverse of a matrix with a special class of directed trees and proved that the conjecture proposed in [15] holds for this special class of directed trees. In this section, we define a new class of simple symmetric directed graphs $\mathbb{D}$ which is an extension of the class considered in [23, Definition 1.2] and showed that the formula obtained in [23, Theorem 1.4] is extendable for the class $\mathbb{D}$ (Theorem 2.8).
Let us now introduce a new class of directed graphs.
Definition 2.1.
Let $\mathbb{D}$ denote the set of all simple symmetric directed graphs $D$ such that each non-pendant vertex of $D$ is adjacent to at least one pendant vertex of $D$.
An important point to note is that the underlying graphs (of this class $\mathbb{D}$) are a special case of cluster networks derived by taking arbitrary graphs as bases and stars as satellites, [3, 30]. Cluster networks are highly relevant in applications in Chemistry since all composite molecules consisting of some amalgamation over a central submolecule can be understood as generalized cluster networks. For instance, they can be used to understand some issues in metal-metal interaction in some molecules, since a cluster network structure can be easily found. In [30], the Kirchhoff index formulae for composite graphs known as join, corona and cluster of two graphs, are presented, in terms of the pieces. The Kirchhoff index formulae and the effective resistances of generalized composite networks, such as generalized cluster or corona network are obtained, in terms of the pieces, in [3]. We refer the reader to [2, 6] for more details on the kirchhoff index of networks.
Example 2.2.
It is clear that the digraph $D_{1}\in\mathbb{D},$ (Fig. 1) while $D_{2}\notin\mathbb{D}$ (Fig. 2). The non-pendant vertex $4$ (in $D_{2}$) is not adjacent to any pendant vertex.
22221111444433337777101010109999888855556666
Figure 1. $D_{1}$
222211114444333377779999888855556666
Figure 2. $D_{2}$
In another way, the property of a digraph $D$ in $\mathbb{D}$ can be written as follows: If $P\subset V(D)$ is the set of pendant vertices, then $\delta(P)=V(D)\backslash P$ or in other words, $P$ is a dense subset of $V(D)$. Sometimes, this property is referred to as the set $P$ dominates $V(D)$. Remember that for a set $F\subset V(D)$ we have $\overline{F}=F\cup\delta(F)$, where $\delta(F)=\{x\in V(D)\backslash F:x\mbox{ is adjacent to }y\mbox{ for some }y\in F\}$. Suppose a directed graph $D$ is in $\mathbb{D}$.
Recall that a directed graph that has a path from each vertex to every other vertex is called strongly connected. A square matrix $A$ is said to be irreducible if $D(A)$ is strongly connected. Let $N(i)$ denote the set of all vertices adjacent to vertex $i$. Suppose a directed graph $D$ is in $\mathbb{D}$. Then, each component of $D$ also belongs to $\mathbb{D}$. Henceforth, we will consider a strongly connected graph in $\mathbb{D}$. Now, we obtain some special properties for the digraphs in $\mathbb{D}$ in the following few results. In general, they are not valid for any simple symmetric digraph.
Proposition 2.3.
Let $D\in\mathbb{D}$ be a strongly connected graph. Then, no non-pendant cycle can belong to a maximum matching of D.
Proof.
In [23, Proposition 2.2], we prove the same result for a tree digraph $D\in\mathbb{D}$. For any digraph $D\in\mathbb{D}$, we can obtain this result by giving the same proof. ∎
Remark 2.4.
Let $D\in\mathbb{D}$ have $k$ non-pendant vertices. Then, the number of 2-cycles in a maximum matching is always $k$. Note that, every non-pendant vertex is matched in any maximum matching of $D$, and both the endpoints of a length three alternating cycle chain are pendant vertices and a length one alternating cycle chain is nothing but a pendant cycle.
Proposition 2.5.
Let $D\in\mathbb{D}$ be a strongly connected graph. We then have the following:
$(i)$
The length of any alternating cycle chain is at most three.
$(ii)$ Any alternating cycle chain between any two vertices is unique.
Proof.
$(i)$ Suppose $D$ has an alternating cycle chain $C$ of length at least five. Then, $C$ must have at least one non-pendant maximum matching cycle, a contradiction to Proposition 2.3.
$(ii)$ Let $C$ be an alternating cycle chain from vertex $i$ to $j$ in $D$. Then, by first part, the length of $C$ is at most three. If the length is one then $C$ is simply a pendant cycle. So, it is unique. Now, let us consider the case when the length of $C$ is three. Since pendant cycles are the only maximum matching cycles (corresponding to any maximum matching), the initial and the terminal vertices of $C$ should be pendant vertices. So, $i$ and $j$ are pendant vertices. Let $\tilde{C}$ be another alternating cycle chain from $i$ to $j$. Since $i$ and $j$ are pendant vertices, the initial and the terminal 2-cycles of $C$ and $\tilde{C}$ must be the same. Here, $\tilde{C}$ is an alternating cycle chain, so its length should be odd. $C$ and $\tilde{C}$ are two different alternating cycle chains with the condition that they have the same initial and terminal 2-cycles. This means that the length of the cycle chain $\tilde{C}$ should be at least five, a contradiction to the fact that the length of any alternating cycle chain in $D$ is at most three (as we proved in part (i)). Therefore, $C$ and $\tilde{C}$ coincide. ∎
Let $A$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Suppose $D(A)$ has $k$ non-pendant vertices. For $i=1,2,\ldots,k$, let $x_{i}$ and $y_{i}$ be (column) vectors of length $r_{i}\in\mathbb{N}$ such that every coordinate is nonzero $(\sum_{i=1}^{k}r_{i}=n-k)$. We shall refer to such vectors as strictly nonzero vectors. Set
$F=\begin{pmatrix}x_{1}^{T}&0&\ldots&0\\ 0&x_{2}^{T}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\ldots&0&x_{k}^{T}\\ \end{pmatrix}\in\mathbb{R}^{k\times(n-k)}$ and $G=\begin{pmatrix}y_{1}&0&\ldots&0\\ 0&y_{2}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\ldots&0&y_{k}\\ \end{pmatrix}\in\mathbb{R}^{(n-k)\times k}$.
Then $A$ can be written permutationally similar to a matrix of the form
$\begin{pmatrix}E&F\\ G&0\\ \end{pmatrix},$ (2.1)
where $E=(e_{ij})\in\mathbb{R}^{k\times k}$. Set $\alpha_{i}=x_{i}^{T}y_{i}$ for all $i\in\{1,2,\ldots,k\}$. Let $W=(W_{ij})$ be the $k\times k$ block matrix, where $W_{ij}=(\frac{e_{ij}}{\alpha_{i}\alpha_{j}})y_{i}x_{j}^{T}$, an $r_{i}\times r_{j}$ matrix. Then, by [22, Theorem 2.5], the group inverse of matrices of the form 2.1 has the $2\times 2$ block form
$\begin{pmatrix}0&~{}Y\\ Z&-W\\ \end{pmatrix},$ (2.2)
where $Y=\frac{x_{1}^{T}}{\alpha_{1}}\oplus\frac{x_{2}^{T}}{\alpha_{2}}\oplus\ldots\oplus\frac{x_{k}^{T}}{\alpha_{k}}$ and $Z=\frac{y_{1}}{\alpha_{1}}\oplus\frac{y_{2}}{\alpha_{2}}\oplus\ldots\oplus\frac{y_{k}}{\alpha_{k}}$.
For a square matrix $A$, $A^{\#}$ may not always exist. For instance, if $A=\begin{pmatrix}0&1\\ 0&0\\ \end{pmatrix},$ then $A^{\#}$ does not exist. In particular, it is easy to show that the group inverse does not exist for any nilpotent matrix. The following proposition provides a necessary and sufficient condition for the existence of the matrix $A^{\#}$ when $D(A)\in\mathbb{D}$, that is the sum of all maximum matchings in $D(A)$ is nonzero.
Proposition 2.6.
Let $A$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Assume that $D(A)$ is strongly connected. Then $A^{\#}$ exists if and only if $\Delta_{A}\neq 0$.
Proof.
$A$ has the form 2.1 since $D(A)\in\mathbb{D}$. Now, set $\alpha_{i}=x_{i}^{T}y_{i}$ for all $i\in\{1,2,\ldots,k\}$. For each $i$, the non-pendant vertex $i$ is adjacent to $r_{i}$ pendant vertices and $\alpha_{i}$ is the sum of cycle products of $r_{i}$ pendant cycles incident $i$. Note that
$FG=\begin{pmatrix}\alpha_{1}&0&\ldots&0\\ 0&\alpha_{2}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\ldots&0&\alpha_{k}\\ \end{pmatrix}\in\mathbb{R}^{k\times k}$ and $F^{\dagger}=\begin{pmatrix}\frac{x_{1}}{x_{1}^{T}x_{1}}&0&\ldots&0\\ 0&\frac{x_{2}}{x_{2}^{T}x_{2}}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\ldots&0&\frac{x_{k}}{x_{k}^{T}x_{k}}\\ \end{pmatrix}$.
By [29, Theorem 3.1], $A^{\#}$ exists if and only if $rank(F)=rank(FG)$. Here, $rank(F)=k$ which implies $A^{\#}$ exists if and only if $\alpha_{i}\neq 0$ for all $i\in\{1,2,\ldots,k\}$. By Proposition 2.3, all the maximum matching cycles are pendant cycles and a maximum matching of $D(A)$ has a set of $k$ pendant cycles incident to $k$ non-pendant vertices (the total number of maximum matching is $\prod_{i=1}^{k}r_{i}$). So, $\Delta_{A}=\prod_{i=1}^{k}\alpha_{i}$ and it follows that $A^{\#}$ exists if and only if $\Delta_{A}\neq 0$. ∎
Let us define more terminology for a strongly connected graph $D(A)\in\mathbb{D}$, where $A$ is a real square matrix. For arbitrary vertices $i$ and $j$ in $D(A)$ denote $\mathbb{M}(i,j)$ to be the set of all maximum matchings $M$ in $D(A)$ such that $C_{m}(i,j)$ is an alternating cycle chain with respect to $M$. Clearly, $\mathbb{M}(i,j)=\mathbb{M}(j,i)$. A necessary condition for the set $\mathbb{M}(i,j)$ to be non-empty is that the length of the path from $i$ to $j$ be odd. If $(i,j,i)$ is a 2-cycle of some maximum matching, then $\mathbb{M}(i,j)$ is non-empty. Two distinct vertices $i$ and $j$ will be called maximally matchable if $\mathbb{M}(i,j)\neq\phi$ and by Proposition 2.5, for a maximally matchable vertices $i,j$, $C_{m}(i,j)$ is unique.
Further, for any maximally matchable vertices $i,j$ and a maximum matching $M\in\mathbb{M}(i,j)$ let $\beta_{\overline{i,j}}(M)$ denote the product of all cycle product, ranging over all the cycles of $M$ that is not contained in the unique cycle chain $C_{m}(i,j)$ in $D(A)$ (product over an empty set is considered to be equal to 1). Since $\mathbb{M}(i,j)=\mathbb{M}(j,i)$, note that $\beta_{\overline{i,j}}(M)=\beta_{\overline{j,i}}(M)$. For a maximum matching $M$ in $D(A)$, $\eta(M)$ denotes the maximum matching product. Set
$\beta_{ij}=\left\{\begin{array}[]{cl}(-1)^{\frac{m-1}{2}}P_{m}(i,j)&\mbox{if }i,j\mbox{ are maximally matchable, }\\ 0&\mbox{if }i,j\mbox{ are not maximally matchable. }\end{array}\right.$
and
$\mu_{ij}=\beta_{ij}\cdot\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M).$ (2.3)
It follows $\mu_{ij}=0$ if $i,j$ are not maximally matchable. This includes the case $i=j$.
In this terminology, a formula for the entries of the group inverse of a matrix $A$ with tree digraph $D(A)\in\mathbb{D}$, proved in [23], is recalled next.
Theorem 2.7.
[23, Theorem 1.4] Let $A$ be an $n\times n$ real matrix with a tree graph $D(A)\in\mathbb{D}$ and assume that $\Delta_{A}\neq 0$. Let $A^{\#}=(\alpha_{ij})$ and let $\mu_{ij}$ be defined as above. Then, $\alpha_{ij}=\frac{\mu_{ij}}{\Delta_{A}}$.
Here is the main result of this section:
Theorem 2.8.
Let $A=(a_{ij})$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Assume that $D(A)$ is a strongly connected graph and $\Delta_{A}\neq 0$. Let $A^{\#}=(\alpha_{ij})$ and let $\mu_{ij}$ be defined in 2.3. Then, $\alpha_{ij}=\frac{\mu_{ij}}{\Delta_{A}}$.
Remark 2.9.
Let $A=(a_{ij})$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Let $\{i_{1},i_{2},\ldots,i_{s+1}\}$ be a set of all pendant vertices with a common neighbour $q$. Then,
$(i)$
$\mathbb{M}=\cup_{m=1}^{s+1}\mathbb{M}(i_{m})$ and $\cap_{m=1}^{s+1}\mathbb{M}(i_{m})=\phi$
$(ii)$ $\sum_{M\in\mathbb{M}(i_{m},q)}\beta_{\overline{i_{m},q}}(M)=\sum_{M\in\mathbb{M}(i_{t},q)}\beta_{\overline{i_{t},q}}(M)$ for all $m,t\in\{1,2,\ldots,s+1\}$
In order to prove the main result, we need three lemmata, which we prove now.
Lemma 2.10.
Let $A=(a_{ij})$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Let $D(A)$ be a strongly connected graph and $\Delta_{A}\neq 0$. Let $B=(b_{ij})$ be the matrix given by $b_{ij}=\frac{\mu_{ij}}{\Delta_{A}},1\leq i,j\leq n$. Then, $AB=BA$.
Proof.
Let $A=(a_{ij})$. Then, $AB=BA$ has the following equivalent form :
$\sum_{k=1}^{n}a_{ik}\mu_{kj}=\sum_{l=1}^{n}\mu_{il}a_{lj}\ \ \ for~{}every\ \ i,j\in\{1,2,\ldots n\}.$ (2.4)
First, we discuss the case $i=j$. Let $\{i_{1},i_{2},\ldots,i_{t}\}\in\{1,2,\ldots,n\}$ be such that for any $m\in\{1,2,\ldots,t\}$, $a_{i,i_{m}}\neq 0$ and the cycle $(i,i_{m},i)$ belongs to some maximum matching in $D(A)$. Since $\mathbb{M}(i,i_{m})=\mathbb{M}(i_{m},i)$ and $\beta_{\overline{i,i_{m}}}(M)=\beta_{\overline{i_{m},i}}(M)$, the expressions on both the sides of equation (2.4) are equal and they equal the common value $\sum_{m=1}^{t}\big{(}a_{ii_{m}}a_{i_{m}i}\sum_{M\in\mathbb{M}(i_{m},i)}\beta_{\overline{i_{m},i}}(M)\big{)}=\sum_{M\in\mathbb{M}(i)}\eta(M)$.
Now, we consider the fact $i\neq j$ and prove 2.4 by considering four cases.
Case (i): $i$ and $j$ are pendant vertices.
Subcase (i.1): $i$ and $j$ have a unique common neighbour. Let $q$ be the common neighbour. Then, $(i,q,i)$ and $(j,q,j)$ are both maximum matching cycles and cannot simultaneously be present in a maximum matching. So, $\{M\backslash\{(i,q,i)\}|~{}M\in\mathbb{M}(i,q)\}=\{M\backslash\{(j,q,j)\}|~{}M\in\mathbb{M}(j,q)\}$. Thus,
$a_{iq}\mu_{qj}$ $=a_{iq}a_{qj}\sum_{M\in\mathbb{M}(q,j)}\beta_{\overline{q,j}}(M)$
$=a_{qj}\Big{(}a_{iq}\sum_{M\in\mathbb{M}(i,q)}\beta_{\overline{i,q}}(M)\Big{)}$
$=\mu_{iq}a_{qj}.$
It follows that equation 2.4 holds.
Subcase (i.2): There is no common neighbour for the vertices $i$ and $j$. Let vertices $i$ and $j$ be adjacent to non-pendant vertices $q$ and $p$, respectively. Then, the length of any cycle chain from $i$ to $j$ is at least three. As we have observed in Remark 2.4, both the end vertices of a length three alternating cycle chain must be pendant vertices, $\mu_{qj}$ and $\mu_{ip}$ are both zero. So, equation 2.4 is vacuously true.
Case (ii): $i$ and $j$ are non-pendant vertices. Since a non-pendant cycle can not be present in a maximum matching and the end vertices of a length three alternating cycle chain are pendant, $\mu_{qj}$ and $\mu_{ip}$ are zero for any arbitrary vertex $q$ adjacent to $i$ and $p$ adjacent to $j$ respectively. So, the expressions on both sides of equation 2.4 are zero.
Case (iii): $i$ is a pendant vertex, while $j$ is a non-pendant vertex. Let $N_{j}=\{j_{1},j_{2},\ldots,j_{s}\}$ be the set of all pendant vertices adjacent to $j$. We consider two subcases.
Subcase (iii.1): $i$ is adjacent to $j$. then, the left hand side of 2.4 is zero and the right hand side equal to
$\sum_{l=1}^{s}\mu_{ij_{l}}a_{j_{l}j}+\sum_{\{m|m\in N(j)\backslash N_{j}\}}\mu_{im}a_{mj}.$ (2.5)
The first sum is zero because $i$ and $j_{l}$ are pendant vertices with common neighbour $j$ for all $l\in\{1,2,\ldots,s\}$. In the second sum, for any $m$, the vertex $m$ is non-pendant and also it is not adjacent to pendant vertex $i$. So, $\mu_{im}$ is zero.
Subcase (iii.2): $i$ is not adjacent to $j$. Let $q$ be the non-pendant vertex adjacent to $i$. The analysis in this case is split into two further subcases, which we consider next.
Subcase (iii.2.a): $q$ is adjacent to $j$. Since non-pendant cycles do not belong to any maximum matching, again, the left hand side of 2.4 is zero, and the right hand side can be split as 2.5. Let $\mathbb{M}_{j_{l}}(i,q)\subseteq\mathbb{M}(i,q)$ be the set of maximum matching containing pendant cycle $(j,j_{l},j)$. Then $\mathbb{M}(i,q)=\cup_{l=1}^{s}\mathbb{M}_{j_{l}}(i,q)$, a mutually disjoint union. Now, the second sum of 2.5 is
$\mu_{iq}a_{qj}$ $=a_{iq}a_{qj}\sum_{M\in\mathbb{M}(i,q)}\beta_{\overline{i,q}}(M)$
$=a_{qj}a_{iq}\sum_{l=1}^{s}\sum_{M\in\mathbb{M}_{j_{l}}(i,q)}\beta_{\overline{i,q}}(M)$
$=a_{qj}a_{iq}\sum_{l=1}^{s}a_{jj_{l}}a_{j_{l}j}\sum_{M\in\mathbb{M}_{j_{l}}(i,q)}\beta_{\overline{i,j_{l}}}(M)$
$=-\sum_{l=1}^{s}a_{j_{l}j}\big{(}-a_{iq}a_{qj}a_{jj_{l}}\big{)}\sum_{M\in\mathbb{M}(i,j_{l})}\beta_{\overline{i,j_{l}}}(M)$
$=-\sum_{l=1}^{s}a_{j_{l}j}\mu_{ij_{l}}.$
So, the right hand side of 2.4 is also zero.
Subcase (iii.2.b): $q$ is not adjacent to $j$. Since $q$ and $j$ both are non-pendant vertices, the left hand side of equation 2.4 is zero. After splitting the right hand side as 2.5, in the first sum, the length of any cycle chain from vertex $i$ to $j_{l}$ must be at least four for any $l\in\{1,2,\ldots,s\}$ and so, $\mu_{ij_{j}}$ is zero. Next, as $m$ is a non-pendant vertex and since it is not adjacent to $i$, the second sum is also zero, showing that the right hand side equals zero.
Case (iv): $i$ is a non-pendant vertex and $j$ is a pendant vertex. Let $N_{i}=\{i_{1},i_{2},\ldots,i_{r}\}$ be the set of all pendant vertices adjacent to $i$. Then, the proof is the same as in Case (iii) by interchanging the roles of $i$ and $j$. ∎
In the next result, we present a graph theoretical interpretation for the entries of the product $AB$, where $A$ and $B$ are as defined in Lemma 2.10.
Corollary 2.11.
Let $A$ and $B$ satisfy the hypotheses of Lemma 2.10. Then,
$(AB)_{ii}=\left\{\begin{array}[]{cl}1&\mbox{if }i\mbox{ is a non-pendant vertex, }\\ \frac{\sum_{M\in\mathbb{M}(i)}\eta(M)}{\Delta_{A}}&\mbox{if }i\mbox{ is a pendant vertex}\end{array}\right.$
while for $i\neq j$,
$(AB)_{ij}=\left\{\begin{array}[]{cl}\frac{a_{qj}\mu_{iq}}{\Delta_{A}}&\mbox{if}i,j\mbox{ are pendant vertices and }\\ &\mbox{have a common neighbour }q,\\ 0&\mbox{otherwise. }\end{array}\right.$
Proof.
It follows from Lemma 2.10 that $(AB)_{ii}=\frac{\sum_{M\in\mathbb{M}(i)}\eta(M)}{\Delta_{A}}$. By Remark 2.4, a non-pendant vertex is matched in every maximum matching, and so for a non-pendant vertex $i$, $(AB)_{ii}=\frac{1}{\Delta_{A}}\cdot\Delta_{A}=1$. For $i\neq j$, the proof is followed by observing case by case in the above Lemma. ∎
Remark 2.12.
In the above corollary, for $i\neq j$, $(AB)_{ij}$ is not equal to zero only when $i,j$ are pendant vertices and have common neighbour.
Lemma 2.13.
Let $A$ and $B$ satisfy the hypotheses of Lemma 2.10. Then, $ABA=A$.
Proof.
To show that $ABA=A$, we show
$\sum_{k=1}^{n}(AB)_{ik}a_{kj}=\left\{\begin{array}[]{ll}a_{ij},&\mbox{when }(i,j)\mbox{ is an edge}\\ 0,&\mbox{when }(i,j)\mbox{ is not an edge}.\end{array}\right.$ (2.6)
Let $c$ be the left hand side of (2.6). Then, $c$ can be written in the form $c=c_{i}+\tilde{c_{i}}$, where
$c_{i}=(AB)_{ii}a_{ij}\ \ \ \mbox{and}\ \ \ \ \tilde{c_{i}}=\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{n}(AB)_{ik}a_{kj}.$
For a pendant vertex $i$, let $\{i_{1},i_{2},\ldots,i_{s}\}$ be the set of all pendant vertices other than $i$, which have a common neighbour.
Case (i): $(i,j)$ is an edge.
Subcase (i.1): $i$ is a non-pendant vertex. By Corollary 2.11, $(AB)_{ii}=1$ and $(AB)_{ik}=0$ for any $k\neq i$. So, $c=a_{ij}$.
Subcase (i.2): $i$ is a pendant vertex. Then, $j$ is a non-pendant vertex adjacent to $i$. By Remark 2.9, $\mathbb{M}=\mathbb{M}(i)\cup_{m=1}^{s}\mathbb{M}(i_{m})$, a mutually disjoint union. By Corollary 2.11, $c_{i}=\frac{\sum_{M\in\mathbb{M}(i)}\eta(M)}{\Delta_{A}}a_{ij}$ and
$\tilde{c_{i}}$ $=\sum_{m=1}^{s}(AB)_{ii_{m}}a_{i_{m}j}$
$=\sum_{m=1}^{s}\Big{(}\frac{a_{ji_{m}}\mu_{ij}}{\Delta_{A}}\Big{)}a_{i_{m}j}$
$=\frac{1}{\Delta_{A}}\sum_{m=1}^{s}a_{ji_{m}}a_{i_{m}j}\Big{(}\beta_{ij}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}$
$=\frac{1}{\Delta_{A}}\sum_{m=1}^{s}a_{ji_{m}}a_{i_{m}j}\Big{(}a_{ij}\sum_{M\in\mathbb{M}(i_{m},j)}\beta_{\overline{i_{m},j}}(M)\Big{)}$
$=\frac{a_{ij}}{\Delta_{A}}\sum_{m=1}^{s}\Big{(}a_{ji_{m}}a_{i_{m}j}\sum_{M\in\mathbb{M}(i_{m},j)}\beta_{\overline{i_{m},j}}(M)\Big{)}$
$=\frac{a_{ij}}{\Delta_{A}}\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m})}\eta(M).$
Thus, $c=\frac{a_{ij}}{\Delta_{A}}\Big{(}\sum_{M\in\mathbb{M}(i)}\eta(M)+\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m})}\eta(M)\Big{)}=\frac{a_{ij}}{\Delta_{A}}\cdot\Delta_{A}=a_{ij}$.
Case (ii): $(i,j)$ is not an edge. Then, $a_{ij}=0$. So, $c=\tilde{c_{i}}$.
Subcase (ii.1): $i$ is a non-pendant vertex. Then, by Remark 2.12, $(AB)_{ik}$ is zero for any $k\neq i$. So, $c=\tilde{c_{i}}=0$.
Subcase (ii.2): $i$ is a pendant vertex. Then, again by Remark 2.12, $\tilde{c_{i}}=\sum_{m=1}^{s}(AB)_{ii_{m}}a_{i_{m}j}$. Since $i$ is not adjacent to $j$, $i_{m}$ is also not adjacent to $j$ for any $m\in\{1,2,\ldots,s\}$. Thus, $c=\tilde{c_{i}}=0$, completing the proof. ∎
Lemma 2.14.
Let $A$ and $B$ satisfy the hypotheses of Lemma 2.10. Then, $BAB=B$.
Proof.
By Lemma 2.10, if we prove $ABB=B$, then we are done. This is equivalent to proving that,
$\sum_{k=1}^{n}(AB)_{ik}\mu_{kj}=\left\{\begin{array}[]{ll}\mu_{ij},&\mbox{if }i,j\mbox{ are maximally matchable, }\\ 0,&\mbox{otherwise}.\end{array}\right.$ (2.7)
Let $b$ be the left hand side of (2.7). Then, $b$ can be written in the form $b=b_{i}+\tilde{b_{i}}$, where
$b_{i}=(AB)_{ii}\mu_{ij}\ \ \ \mbox{and}\ \ \ \ \tilde{b_{i}}=\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{n}(AB)_{ik}\mu_{kj}.$
Case (i): $i,j$ are maximally matchable. Then, by Proposition 2.5, the length of alternating cycle chain from $i$ to $j$ is at most three.
Subcase (i.1): The length of alternating cycle chain from $i$ to $j$ is three. Then, $i$ and $j$ are both must be pendant vertex. Let $\{i_{1},i_{2},\ldots,i_{s}\}$ be the set of all pendant vertices other than $i$, which have a common neighbour $q$. Let the unique alternating cycle chain from $i$ to $j$ be $((i,q,i),(q,z,q),(z,j,z))$. Again, by Corollary 2.11, $b_{i}=\frac{\sum_{M\in\mathbb{M}(i)}\eta(M)}{\Delta_{A}}\mu_{ij}$ and
$\tilde{b_{i}}$ $=\sum_{m=1}^{s}(AB)_{ii_{m}}\mu_{i_{m}j}$
$=\sum_{m=1}^{s}\Big{(}\frac{a_{qi_{m}}\mu_{iq}}{\Delta_{A}}\Big{)}\mu_{i_{m}j}$
$=\frac{1}{\Delta_{A}}\sum_{m=1}^{s}a_{qi_{m}}\Big{(}\beta_{iq}\sum_{M\in\mathbb{M}(i,q)}\beta_{\overline{i,q}}(M)\Big{)}\Big{(}\beta_{i_{m}j}\sum_{M\in\mathbb{M}(i_{m},j)}\beta_{\overline{i_{m},j}}(M)\Big{)}$
$=\frac{1}{\Delta_{A}}\sum_{m=1}^{s}a_{qi_{m}}\Big{(}a_{iq}\sum_{M\in\mathbb{M}(i_{m},q)}\beta_{\overline{i_{m},q}}(M)\Big{)}\Big{(}-a_{i_{m}q}a_{qz}a_{zj}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}$
$=\frac{1}{\Delta_{A}}\sum_{m=1}^{s}\Big{(}a_{i_{m}q}a_{qi_{m}}\sum_{M\in\mathbb{M}(i_{m},q)}\beta_{\overline{i_{m},q}}(M)\Big{)}\Big{(}-a_{iq}a_{qz}a_{zj}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}$
$=\frac{1}{\Delta_{A}}\sum_{m=1}^{s}\Big{(}\sum_{M\in\mathbb{M}(i_{m})}\eta(M)\Big{)}\Big{(}\beta_{ij}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}$
$=\frac{\mu_{ij}}{\Delta_{A}}\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m})}\eta(M).$
Thus, by Remark 2.9, $b=\frac{\mu_{ij}}{\Delta_{A}}\Big{(}\sum_{M\in\mathbb{M}(i)}\eta(M)+\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m})}\eta(M)\Big{)}=\frac{\mu_{ij}}{\Delta_{A}}\cdot\Delta_{A}=\mu_{ij}$.
Subcase (i.2): The length of alternating cycle chain from $i$ to $j$ is one. Then, $i$ is adjacent to $j$.
Subcase (i.2.a): $i$ is a non-pendant vertex. Then, $(AB)_{ii}=1$ and $(AB)_{ik}=0$ for any $k\neq i$. So, $b=\mu_{ij}$.
Subcase (i.2.b): $i$ is a pendant vertex. Let $\{i_{1},i_{2},\ldots,i_{s}\}$ be the set of all pendant vertices other than $i$, which have a common neighbour $j$. Then, by Remark 2.9 and 2.12, $b_{i}=\frac{\sum_{M\in\mathbb{M}(i)}\eta(M)}{\Delta_{A}}\mu_{ij}$ and
$\tilde{b_{i}}$ $=\sum_{m=1}^{s}(AB)_{ii_{m}}\mu_{i_{m}j}$
$=\sum_{m=1}^{s}\Big{(}\frac{a_{ji_{m}}\mu_{ij}}{\Delta_{A}}\Big{)}\mu_{i_{m}j}$
$=\frac{\mu_{ij}}{\Delta_{A}}\sum_{m=1}^{s}a_{ji_{m}}\Big{(}\beta_{i_{m}j}\sum_{M\in\mathbb{M}(i_{m},j)}\beta_{\overline{i_{m},j}}(M)\Big{)}$
$=\frac{\mu_{ij}}{\Delta_{A}}\sum_{m=1}^{s}\Big{(}a_{ji_{m}}a_{i_{m}j}\sum_{M\in\mathbb{M}(i_{m},j)}\beta_{\overline{i_{m},j}}(M)\Big{)}$
$=\frac{\mu_{ij}}{\Delta_{A}}\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m})}\eta(M).$
Thus, by Remark 2.9, $b=\frac{\mu_{ij}}{\Delta_{A}}\Big{(}\sum_{M\in\mathbb{M}(i)}\eta(M)+\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m})}\eta(M)\Big{)}=\frac{\mu_{ij}}{\Delta_{A}}\cdot\Delta_{A}=\mu_{ij}$.
Case (ii): $i,j$ are not maximally matchable. Then, $\mu_{ij}=0$ and $b=\tilde{b_{i}}$.
Subcase (ii.1): $i$ is pendant vertex. Let $\{i_{1},i_{2},\ldots,i_{s}\}$ be the set of all pendant vertices other than $i$, which have a common neighbour. Then, by Remark 2.12, $\tilde{b_{i}}=\sum_{m=1}^{s}(AB)_{ii_{m}}\mu_{i_{m}j}$. Since $\{i,i_{1},i_{2},\ldots,i_{s}\}$ be a set of all pendant vertices with a common neighbour and $i,j$ are not maximally matchable, $i_{m},j$ are also not maximally matchable for any $m\in\{1,2,\ldots,s\}$. So, $\mu_{i_{m}j}=0$ for all $m\in\{1,2,\ldots,s\}$. Thus, $b=\tilde{b_{i}}=0$.
Subcase (ii.2): $i$ is non-pendant vertex. Then, again by Remark 2.12, $(AB)_{ik}=0$ for any $k\neq i$. Thus, $b=\tilde{b_{i}}=0$, completing the proof. ∎
Here is an illustration.
Example 2.15.
Consider the matrix
$A=\begin{pmatrix}~{}0&-2&~{}0&~{}2&2&1&~{}0&0&0&0\\ -1&~{}0&~{}1&~{}0&0&0&-3&0&0&0\\ ~{}0&~{}3&~{}0&-1&0&0&~{}0&2&2&0\\ -1&~{}0&~{}1&~{}0&0&0&~{}0&0&0&2\\ ~{}1&~{}0&~{}0&~{}0&0&0&~{}0&0&0&0\\ -3&~{}0&~{}0&~{}0&0&0&~{}0&0&0&0\\ ~{}0&-2&~{}0&~{}0&0&0&~{}0&0&0&0\\ ~{}0&~{}0&~{}3&~{}0&0&0&~{}0&0&0&0\\ ~{}0&~{}0&-1&~{}0&0&0&~{}0&0&0&0\\ ~{}0&~{}0&~{}0&~{}2&0&0&~{}0&0&0&0\\ \end{pmatrix}.$
Then, $D(A)$ is the digraph $D_{1}$ in Figure 1. Here, all the maximum matchings of $D(A)$ are given by:
$M_{1}=\{(1,5,1),(2,7,2),(3,8,3),(4,10,4)\}$
$M_{2}=\{(1,5,1),(2,7,2),(3,9,3),(4,10,4)\}$
$M_{3}=\{(1,6,1),(2,7,2),(3,8,3),(4,10,4)\}$
$M_{4}=\{(1,6,1),(2,7,2),(3,9,3),(4,10,4)\}$
So, $\Delta_{A}=288+(-96)+(-432)+144=-96$. Let $A^{\#}=(\alpha_{ij})$. Let us compute $\alpha_{57}$. First, $P_{3}(5,7)=1\times(-2)\times(-3)=6$. Note that $C_{3}(5,7)$ cycle chain is alternating with respect to the maximum matchings $M_{1}$ and $M_{2}$. Thus, $\beta_{57}=(-1)\times 6=-6$, $\beta_{\overline{5,7}}(M_{1})=6\times 4=24$ and $\beta_{\overline{5,7}}(M_{2})=(-2)\times 4=-8$. So,
$\mu_{57}=(-6)\times(24-8)=-96.$
Therefore, $\alpha_{57}=\frac{-96}{-96}=1$.
3. Necessary and Sufficient Condition of $A$ such that $D(A),D(A^{\#})\in\mathbb{D}$
Before we proceed, let us provide two examples. First we consider a matrix
$A=\begin{pmatrix}0&~{}2&-1&0&0\\ 1&~{}0&~{}0&1&1\\ 1&~{}0&~{}0&0&0\\ 0&-2&~{}0&0&0\\ 0&-2&~{}0&0&0\\ \end{pmatrix}.$
Then $D(A)\in\mathbb{D}$ and
$A^{\#}=\begin{pmatrix}~{}0&0&~{}1&~{}0&~{}0\\ ~{}0&0&~{}0&-\frac{1}{4}&-\frac{1}{4}\\ -1&0&~{}0&-\frac{1}{2}&-\frac{1}{2}\\ ~{}0&\frac{1}{2}&-\frac{1}{2}&~{}0&~{}0\\ ~{}0&\frac{1}{2}&-\frac{1}{2}&~{}0&~{}0\\ \end{pmatrix}.$
Since the non-pendant vertices 2, 4, 5 are not adjacent to any pendant vertex in $D(A^{\#})$, $D(A^{\#})\notin\mathbb{D}$. On the other hand, for a matrix
$B=\begin{pmatrix}~{}0&1&1&2&-1\\ -1&0&0&0&~{}0\\ ~{}2&0&0&0&~{}0\\ ~{}1&0&0&0&~{}0\\ ~{}2&0&0&0&~{}0\\ \end{pmatrix},$
$D(B)$ is a star tree and belongs to $\mathbb{D}$. Since $B^{\#}=B$, $D(B^{\#})$ also belongs to $\mathbb{D}$. The two examples above serve a good motivation for our main result (Theorem 3.5) in this section.
The first result in this section shows that the $ij$-th entry of the group inverse of a real matrix $A$ is non-zero when $i,j$ are maximally matchable in $D(A)\in\mathbb{D}$.
Corollary 3.1.
Let $A$ satisfy the hypothesis of Theorem 2.8. If $i,j$ are maximally matchable and $A^{\#}=(\alpha_{ij})$, then $\alpha_{ij}\neq 0$.
Proof.
By Proposition 2.5, the length of the alternating cycle chain from $i$ to $j$ is at most three, since $D(A)\in\mathbb{D}$.
Case (i): The length of the alternating cycle chain is one, so that $(i,j,i)$ is a pendant cycle. Without loss of generality, let $i$ be the pendant vertex. Let $\{i_{1}(=i),i_{2},\ldots,i_{s}\}$ be the set of pendant vertices which have $j$, the non-pendant vertex, as a common neighbour. Then, $\mathbb{M}=\cup_{m=1}^{s}\mathbb{M}(i_{m},j),$ a disjoint union. Now,
$\Delta_{A}$ $=\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m},j)}\eta(M)$
$=\sum_{m=1}^{s}a_{i_{m}j}a_{ji_{m}}\sum_{M\in\mathbb{M}(i_{m},j)}\beta_{\overline{i_{m},j}}(M)$
$=\sum_{m=1}^{s}a_{i_{m}j}a_{ji_{m}}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)$
$=\Big{(}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}\sum_{m=1}^{s}a_{i_{m}j}a_{ji_{m}}.$
Since $\Delta_{A}\neq 0$, $\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\neq 0$ and so, $\alpha_{ij}=\frac{\beta_{ij}}{\Delta_{A}}\Big{(}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}\neq 0$.
Case (ii): The length of the alternating cycle chain is three. Let the unique alternating cycle chain from $i$ to $j$ be $((i,q,i),(q,z,q),(z,j,z))$. So, $i,j$ are pendant while $q,z$ are non-pendant vertices. Let $\{i_{1}(=i),i_{2},\ldots,i_{s}\}$ and $\{j_{1}(=j),j_{2},\ldots,j_{p}\}$ be the set of all pendant vertices having $q$ and $p$ as a common neighbour, respectively. Then, $\mathbb{M}=\cup_{m=1}^{s}\cup_{l=1}^{p}\mathbb{M}(i_{m},j_{l})$, again a mutually disjoint union. Now, we obtain
$\Delta_{A}$ $=\sum_{m=1}^{s}\sum_{l=1}^{p}\sum_{M\in\mathbb{M}(i_{m},j_{l})}\eta(M)$
$=\sum_{m=1}^{s}\sum_{l=1}^{p}(a_{i_{m}q}a_{qi_{m}})(a_{j_{l}z}a_{zj_{l}})\sum_{M\in\mathbb{M}(i_{m},j_{l})}\beta_{\overline{i_{m},j_{l}}}(M)$
$=\sum_{m=1}^{s}\sum_{l=1}^{p}(a_{i_{m}q}a_{qi_{m}})(a_{j_{l}z}a_{zj_{l}})\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)$
$=\Big{(}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}\sum_{m=1}^{s}\sum_{l=1}^{p}(a_{i_{m}q}a_{qi_{m}})(a_{j_{l}z}a_{zj_{l}}).$
Again, since $\Delta_{A}\neq 0$, $\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\neq 0$ and so, $\alpha_{ij}\neq 0$. ∎
In general, it is not true that the digraph corresponding to the group inverse of a matrix with a simple symmetric digraph is again a simple symmetric digraph, shown by the following example.
Example 3.2.
Consider a matrix with simple symmetric digraph,
[…]
It is clear that $D(A)$ is a simple symmetric digraph, while $D(A^{\#})$ is not a simple symmetric digraph.
In the following result, we show that when $D(A)\in\mathbb{D}$ for a real matrix $A$, $D(A^{\#})$ is a simple symmetric digraph.
Corollary 3.3.
Let $A$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Assume that $D(A)$ is strongly connected and $\Delta_{A}\neq 0$. Then, $D(A^{\#})$ is a strongly connected simple symmetric directed graph.
Proof.
Since $D(A)$ is strongly connected, $A$ is an irreducible matrix. Then by [18, Lemma 2.4], $A^{\#}$ is irreducible. So, $D(A^{\#})$ is strongly connected. Let $A^{\#}=(\alpha_{ij})$. Then $\mu_{ii}=0$ implies $\alpha_{ii}=0$ and so, $D(A^{\#})$ is simple. For any $i$ and $j$, suppose $\alpha_{ij}\neq 0$. Then, by Theorem 2.8, $i,j$ are maximally matchable. Clearly, $j,i$ are also maximally matchable and so, from Corollary 3.1, $\alpha_{ji}\neq 0$. Thus, $A^{\#}$ is a combinatorially symmetric matrix and $D(A^{\#})$ is a symmetric directed graph. ∎
Recall that a simple symmetric directed graph is said to be corona graph if each non-pendant vertex is adjacent to exactly one pendant vertex.
Proposition 3.4.
Let $A=(a_{ij})$ be an $n\times n$ real matrix such that $\Delta_{A}\neq 0$.
$(i)$
If $D(A)$ is a star tree graph, then $D(A^{\#})$ is a star tree graph.
$(ii)$ If $D(A)$ is a corona graph, then $D(A^{\#})$ is a corona graph.
Proof.
$(i)$ Let the vertex set of $D(A)$ be $\{1,2,\ldots,n\}$. Without loss of generality, let the center vertex of $D(A)$ be $n$. Then, the only non zero entries in $A$ are $a_{in},a_{ni}$ for all $11\leq i\leq n-$ 1. On the other hand, the maximum matchings of $D(A)$ are $M_{i}=\{(n,i,n)\}$ and so, the only maximum matchable pair of vertices are $n,i$ for all $11\leq i\leq n-$ 1. Let $A^{\#}=(\alpha_{ij})$. Then, by Theorem 2.8 and Corollary 3.1, the only non zero entries of $A^{\#}$ are $\alpha_{in},\alpha_{ni}$ for all $11\leq i\leq n-$ 1. Thus, $D(A^{\#})$ is a star tree graph.
$(ii)$ If $D(A)$ is a corona graph, then $n$ is even. Let the vertex set of $D(A)$ be $\{1,2,\ldots,\frac{n}{2},\frac{n}{2}+1,\ldots n\}$. Without loss of generality, let the set of non-pendant vertices and pendant vertices of $D(A)$ be $\{1,2,\ldots,\frac{n}{2}\}$ and $\{\frac{n}{2}+1,\frac{n}{2}+2,\ldots n\}$, respectively, where vertex $k$ is adjacent to vertex $\frac{n}{2}+k$ for all $1\leq k\leq\frac{n}{2}$. Then, the matrix $A$ is of the form 2.1, where $F$ and $G$ both are diagonal matrices of order $\frac{n}{2}$ with non-zero entries in the diagonal. Also, $A$ is invertible and
$A^{\#}=A^{-1}=\begin{pmatrix}0&G^{-1}\\ F^{-1}&-F^{-1}EG^{-1}\\ \end{pmatrix}.$
Since $F$ and $G$ are both diagonal matrices with non-zero entries in the diagonal, the zero non-zero pattern of $E$ and $-F^{-1}EG^{-1}$ will be the same. Then, there exists a permutation matrix $P$ such that
$P^{-1}A^{\#}P=\begin{pmatrix}-F^{-1}EG^{-1}&F^{-1}\\ G^{-1}&0\\ \end{pmatrix}.$
Since the zero non-zero pattern of $A$ and $P^{-1}A^{\#}P$ are same, $D(A^{\#})$ is also a corona graph. ∎
Finally, we classify all the real square matrix $A$ such that $D(A)$ and $D(A^{\#})$ both are in the class $\mathbb{D}$, that is $D(A)$ is either a corona graph or a star graph.
Theorem 3.5.
Let $A$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Assume that $D(A)$ is strongly connected and $\Delta_{A}\neq 0$. Then $D(A^{\#})\in\mathbb{D}$ if and only if $D(A)$ is either a corona graph or a star tree graph.
Proof.
Let $D(A)$ be a corona or star tree graph. Then, by Proposition 3.4, $D(A^{\#})\in\mathbb{D}$. Conversely, suppose $D(A)$ is neither a corona nor a star tree graph. Since $D(A)\in\mathbb{D}$, there exists a non-pendant vertex adjacent to more than one pendant vertices and at least two non-pendant vertices in $D(A)$. Let $i$ be a non-pendant vertex adjacent to $s\geq 2$ pendant vertices $\{i_{1},i_{2},\ldots,i_{s}\}$ and $i$ is adjacent to another non-pendant vertex $j$. The alternating paths that start or end with vertex $i$ are $((i,i_{m},i))$ for all $1\leq m\leq s$. So, $i$ is adjacent to exactly $s$ vertices $\{i_{1},i_{2},\ldots,i_{s}\}$ in $D(A^{\#})$. Let $j$ be adjacent to a pendant vertex $\tilde{j}$. Then, for any $1\leq m\leq s$, $((i_{m},i,i_{m}))$ and $((i_{m},i,i_{m}),(i,j,i),(j,\tilde{j},j))$ are two alternating paths start from vertex $i_{m}$. So, vertex $i_{m}$ is non-pendant in $D(A^{\#})$ for all $1\leq m\leq s$. Since we have a non-pendant vertex $i$ which is not adjacent to any pendant vertices in $D(A^{\#})$, $D(A^{\#})$ does not belong to $\mathbb{D}$. ∎
Acknowledgments
The author was supported by Indian Institute of Technology Gandhinagar Post-Doctoral Fellowship IP/IITGN/MATH/PC/2223/20.